OpenStudy (anonymous):

how many order (a,b) such that : (a^3 + 1)/(ab - 1) is an integer with a, b are integer number

OpenStudy (anonymous):

@rational

OpenStudy (dan815):

oh honey

OpenStudy (dan815):

do cube root factoring first

OpenStudy (anonymous):

a^3 + 1 = (a + 1)(a^2 - a + 1)

OpenStudy (dan815):

okay

OpenStudy (dan815):

(a + 1)(a^2 - a + 1) ----------------- (ab-1)

OpenStudy (dan815):

multiply by conjugate see where that gets us

OpenStudy (dan815):

or well lets consider the bounds here

OpenStudy (anonymous):

conyugate of ab - 1 ?

OpenStudy (dan815):

ya nvm lets do conjugate

OpenStudy (dan815):

can u tell me some pretext to this problem like are u working on mod functions or remainder theorem or anything?

OpenStudy (dan815):

(a + 1)(a^2 - a + 1) ----------------- = K , where K is an integer (ab-1)

OpenStudy (anonymous):

this is just theory number question from my book but idont know what the start to do it

OpenStudy (dan815):

maybe this is obvious but b=a^2+2/a , is one form of solution right

OpenStudy (dan815):

and b=2/a

OpenStudy (anonymous):

how you get that ?

OpenStudy (dan815):

by look at ab-1

OpenStudy (dan815):

since the top has (a+1) and its (a^3+1)

OpenStudy (dan815):

also from the factor (a^2-a+1) we get that ab-1=a^2-a+1 b=(a^2-a+2)/a b=a-1+2/a

OpenStudy (dan815):

getting the int solutions for these will give us some solutions but i dont think these wud be all of the solutions

OpenStudy (dan815):

from b=2/a (1,2) is the only solution u can get from here

OpenStudy (dan815):

also (2,1) should work as well since that also makes the denominator 1

OpenStudy (dan815):

anything divided by 1 will be an integer so we are fine there

OpenStudy (dan815):

any int* div by 1

OpenStudy (anonymous):

wait, how to get b = 2/a ?

OpenStudy (dan815):

that makes the denominator 1

OpenStudy (dan815):

ab-1 =1 b=2/a

OpenStudy (dan815):

another way to get 2/a is that

OpenStudy (anonymous):

oh, okay :D

OpenStudy (dan815):

(a^3 + 1) look at its factors (a^3 + 1)=(a + 1)(a^2 - a + 1) so if ab-1 = a+1 or ab-1=a^2-a+1

OpenStudy (dan815):

we know it will divide with the result as an integer if this holds true

OpenStudy (dan815):

so u get ab-1=a+1 b=(a+2)/a =1+2/a again this leads us back to the fact that b=1+2/a so integer is coming from the fact that a can be 1 and 2 here

OpenStudy (dan815):

all other integers of a will give us a remainder

OpenStudy (dan815):

theresore we have a new solution here

OpenStudy (dan815):

(1,2) (2,1) (2,3) (2,2)

OpenStudy (dan815):

lets see if theres anymore from the other equations

OpenStudy (anonymous):

for b = 1 + 2/a than a = {+-1 , +-2} ?

OpenStudy (dan815):

sorry that shud be (1,3) not (2,3)

OpenStudy (dan815):

oh yes negative too good point

OpenStudy (dan815):

oh and there is a restriction which is ab=/=1

OpenStudy (dan815):

(1,1) is basically the only way that will happen and its not a solution anyway

OpenStudy (dan815):

and -1,-1

OpenStudy (dan815):

also from the factor (a^2-a+1) we get that ab-1=a^2-a+1 b=(a^2-a+2)/a b=a-1+2/a

OpenStudy (anonymous):

from case b = 1 + 2/a i get (a,b) = (1,2),(-1,-1),(2,2),(2,0)

OpenStudy (dan815):

tell me the extra solutions from that equation now

OpenStudy (dan815):

ok so we throw out (-1,-1)

OpenStudy (dan815):

like we agreed

OpenStudy (dan815):

u are missing solution there

OpenStudy (dan815):

(1,3)

OpenStudy (dan815):

nvm just that,

OpenStudy (dan815):

b=a-1+2/a a=1 b=2 a=-1 b=-2-2 = -4 a=2 and -2 b=2, -4

OpenStudy (dan815):

1,2 -1,-4 2,2 -2,-4

OpenStudy (anonymous):

ok, for the solution i can cover it. let's ta2e conclusion as all cases satisfy : * ab -1 = a + 1 ** ab - 1 = a^2 - a + 1 only 2 cases right ?

OpenStudy (anonymous):

*take

OpenStudy (dan815):

there more

OpenStudy (dan815):

i got my cases from this (a^3+1) =ab-1 so b=a^2+2/a ------------------- (a+1)=ab-1 b=1+2/a ------------------ ab - 1 = a^2 - a + 1 b=a-1+2/a ------------------- and ab-1=1 b=2/a

OpenStudy (dan815):

get solutions from those 4

OpenStudy (dan815):

not sure if there anything im missing

OpenStudy (anonymous):

ok, nove . i will cover all cases to get all solutions. thank you very much for your explain.

OpenStudy (anonymous):

*dan815 how about ab - 1 = -1 ? is this make extra case ?

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):

oh yes! it definately does

OpenStudy (anonymous):

integer number divided by -1 = integer also

OpenStudy (anonymous):

thank you :D

OpenStudy (dan815):

well actually it might be covered though

OpenStudy (dan815):

did we see that 0,0 is a solution anywhere

OpenStudy (anonymous):

yes from the case ab - 1 = -1 make ab = 0 (a,b) = (0,0)

OpenStudy (dan815):

oh true wow there an infinite case u found!

OpenStudy (dan815):

(N,0), where N is any integer

OpenStudy (dan815):

(0,N) , Where N is any integer another infinite case solution

OpenStudy (anonymous):

wait, ab = 0 solution a = 0 or b = 0

OpenStudy (dan815):

and a=0 and b= anything a=anythign and b=0

OpenStudy (anonymous):

so, it would be infinity order for (a,b) ?

OpenStudy (dan815):

yep!

OpenStudy (anonymous):

wah... :D